

********************************************************************
UARS WEEKLY STATUS REPORT - 09 April 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
68% of each day's data amounting to about 500 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset, and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1996039.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 March 2004 (Orbit 68623, GMT Day 087) through
02 April 2004 (Orbit 68728, GMT Day 093).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 087/2001 68635 POWER OFF
SUSIM 087/2112 68636 POWER OFF
SOLSTICE 087/2219 68637 TO DAY-ONLY OPS
SOLSTICE 091/1354 68692 DAY-ONLY TO FULL-TIME
SUSIM 091/1717 68694 POWER ON
PEM 093/1348 68722 POWER ON ZEPS/HEPS DAY-only
A Reverse-to-Forward Yaw Around maneuver (number 128) was successfully
executed in Orbit 68662 beginning 089/14:51:09Z and completing at
089/15:28:19Z on 29 March 2004. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 55.43 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.08 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.74 and 20.27 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 07
May 2004 (DOY 128).
The beta angle ranged between the angles of 11.5 degrees to 18.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 68 degrees on 17 April 2004 (DOY 108).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 1.27 degrees C. Additional
battery performance data for each battery on 26 March and 02 April
2004 are provided in Appendix D.
The clock error ranged from 10.2 to -9.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
PEM = 02
SUSIM = 02
SOLSTICE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT transitioned to the 3-TDRS mode of operations on 17 March
2004. The FOT is now taking 35-38 events per day (up from 26-28 per
day under 2-TDRS operations) and collecting data for an additional 18-
20% more of each orbit. So far no problems have been found with this
mode and all subsystems remain nominal.
MEETINGS
The FOT attended a meeting with the HRDI team on 02 April 2004 in
order to follow up on the attitude error issue. Data sets are still
be collected and analyzed. The FOT will meet with the HRDI team again
in two weeks to discuss their findings.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 07 May 2004 (DOY 128). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 April 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68626 087/0542 04-093 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68635 087/2001 04-094 PEM POWER OFF
68636 087/2112 04-095 SUSIM POWER OFF
68637 087/2219 04-096 SOLSTICE TO DAY-ONLY OPS
68647 088/1456 04-097 MACS YAW VT-6
68662 089/1416 04-097 MACS YAW
68682 090/2155 04-098 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68692 091/1354 04-099 SOLSTICE DAY-ONLY TO FULL-TIME
68693 091/1616 04-097 MACS YAW VT-5
68694 091/1717 04-100 SUSIM POWER ON (PART 1)
68706 092/1309 04-100 SUSIM POWER ON (PART 2)
68700 092/0332 04-101 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68722 093/1348 04-102 PEM POWER ON ZEPS/HEPS DAY-ONLY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 April 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 April 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68626 087/0546 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68682 090/2155 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68700 092/0334 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 April 2004
26 March 2004 - GMT Day 086 Beta = 12.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-134.4 +5.79/+3.87 +28.0/-14.4 23.7 1.02 19.8
3 +22.4/-089.6 +5.14/+2.60 +27.2/-14.4 23.5 1.02 19.5
02 April 2004 - GMT Day 093 Beta = 18.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-22.4 +4.50/+3.23 +28.4/-12.8 26.1 1.03 19.3
3 +22.4/-33.6 +3.23/+2.28 +27.2/-12.4 26.1 1.01 19.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
68647 088/1456 LV5 -> LV6
68693 091/1616 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 April 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-016 FOT 68708 092/1613 35490 Drop out due to HW @ WSGT
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT See List Mar 04 30832 Generic Late Acquisition
04-016 FOT 68708 092/1613 35490 Drop out due to HW @ WSGT